can u plz help me

i know answer 1 is 5.58 m/s, but i am looking for answer 2 please
Answer #1

distance = 6.7 + 1.2 * 12 = 21.1 m
Answer #2

1) using d= ut
6.7 = 1.2 x u
u= 5.58 m/s


2) distance = 6.7 + (12 x 1.2) = 21.1 m